# Non-Relativistic Hyperboloidal Method for Quasinormal Modes

### v1

@@ Appendix A Appendix
A.1 1. Equation of Motion for the $\hat{\psi}$ Field
@@
$J_{1}$ $=v_{a}^{2}\Big{[}V+\Delta(1-x^{2})-\Delta^{2}x^{2}+2(1+\Delta)x(1-x^{2})\partial_{x}-(1-x^{2})^{2}\partial_{x}^{2}\Big{]},$
$J_{2}$ $=v_{a}\Big{[}iv_{a}+1-x^{2}-2\Delta x^{2}+2x(1-x^{2})\partial_{x}\Big{]}.$ (A1)
@@
$L_{1}^{a}$ $=v_{a}^{2}\Big{[}\tilde{V}-(1+\Delta)\Delta-2(1+\Delta)x\partial_{x}+(2-x^{2})\partial_{x}^{2}\Big{]},$ $L_{2}^{a}=v_{a}\Big{[}{-}1-2\Delta-2x\partial_{x}\Big{]},$
$J_{1}^{b}$ $=v_{a}^{2}\Big{[}V_{0}+xV_{1}+\Delta+2(1+\Delta)x\partial_{x}-\partial_{x}^{2}\Big{]},$ $J_{2}^{b}=v_{a}\Big{[}iv_{a}+1+2x\partial_{x}\Big{]},\hskip 8.0pt$ (A3)
@@
A.2 2. Determinant of the Matrix $L^{N}-\Omega\text{Id}$
@@
The QNM spectrum is given when this determinant vanishes, meaning we can scale $\Omega^{2}\text{Id}-i\Omega L_{2}^{N}-L_{1}^{N}$ without changing our result. First, we substitute for the discretized operators $L_{1}^{N}$ and $L_{2}^{N}$, making use of the asymptotic dispersion relation to rewrite $v_{a}$ and $v_{p}$ in terms of the real and imaginary parts of $\Omega$. Then, after an overall scaling of the matrix, we arrive at the condition $\det(M)=0$, with

### v4

@@ Appendix A Appendix
A.1 1. Boundedness of QNMs on the Spatial Boundaries
The new coordinates of Eqs. (2) and (3) are a simple modification of Bizoń-Mach coordinates, parametrized with the asymptotic velocity, $v_{g}$. The explicit coordinate transformation may be written as
$\tau=t-(v_{g})^{-1}\log\cosh(r),\quad x=\tanh(r),$ (A1)
If we consider the asymptotic behaviour of these coordinate functions as $r\to\pm\infty$, then we find that
$\tau\to t\pm(v_{g})^{-1}(r+\log(2)),$ (A2)
In the original coordinates, this reads
$\tau\to(v_{g})^{-1}(v_{g}t\pm r)\pm(v_{g})^{-1}\log(2),$ (A3)
from which we see that contours of constant $\tau$ correspond to trajectories propagating outwards with velocity, $v_{g}$.
The asymptotic form of the QNM solutions may be written
$\phi\sim\exp(-\text{Im}(K)r+\text{Im}(\Omega)t)\exp(i\text{Re}(K)r-i\text{Re}(\Omega)t),$ (A4)
where we have separated out evolution of the amplitude and phase. For complex frequencies and wavenumbers, we generalize the group velocity, given by $\text{Im}(\Omega)/\text{Im}(K)$, to correspond to the transport of amplitude.
The boundedness of QNM solutions on the boundaries of the compactified space follows when we consider QNMs whose group velocities at the right and left boundaries are $v_{g}$ and $-v_{g}$, respectively. The contours of constant $\tau$ define the compactified space, and it follows from Eqs. (A3) and (A4) that the asymptotic evolution of amplitude along these contours vanishes for QNMs with the stated group velocities. As a result, these QNMs take on definite amplitudes at the boundaries of the compactified space, $x=\pm 1$, and are bounded everywhere on the space.
A.2 2. Relationship between $\Omega$ and the Asymptotic Velocities
The asymptotic dispersion relation associated with Eq. (1) is $\Omega=-K^{2}$. The frequency and wavenumber are both complex quantities in the case of QNMs, and we begin by separating real from imaginary parts, to obtain
$\text{Re}(\Omega)=(\text{Im}(K)^{2}-\text{Re}(K)^{2}),\quad\text{Im}(\Omega)=-2\text{Re}(K)\text{Im}(K).$ (A5)
Inverting these to obtain the real and imaginary parts of the wavenumber, we arrive at
$\text{Re}(K)=\pm\sqrt{\frac{1}{2}(|\Omega|-\text{Re}(\Omega))},\quad\text{Im}(K)=\pm\sqrt{\frac{1}{2}(|\Omega|+\text{Re}(\Omega))},$ (A6)
where the sign choices are correlated. As discussed above, Eq. (A4) shows that the asymptotic group velocity is the ratio of the imaginary parts of the frequency and wavenumber. Likewise, it shows that the asymptotic phase velocity is the ratio of their real parts. Identifying the parameters $v_{g}$ and $v_{p}$ with the group and phase velocities at the right spatial boundary, $x=1$, we obtain
$v_{p}=\frac{\text{Re}(\Omega)}{\text{Re}(K_{+})}=-\frac{\sqrt{2}\text{Re}(\Omega)}{\sqrt{|\Omega|-\text{Re}(\Omega)}},\quad v_{g}=\frac{\text{Im}(\Omega)}{\text{Im}(K_{+})}=-\frac{\sqrt{2}\text{Im}(\Omega)}{\sqrt{|\Omega|+\text{Re}(\Omega)}}.$ (A7)
These give rise to a pair of equations,
$2\text{Re}(\Omega)^{2}-v_{p}^{2}|\Omega|+v_{p}^{2}\text{Re}(\Omega)=0,\quad 2\text{Im}(\Omega)^{2}-v_{g}^{2}|\Omega|-v_{g}^{2}\text{Re}(\Omega)=0,$ (A8)
which relate the real and imaginary parts of $\Omega$ to the real parameters, $v_{g}$ and $v_{p}$. Requiring QNM solutions to decay in time and have outgoing group velocities at the boundaries, we can solve for $\Omega$ and obtain Eq. (5).
A.3 3. Equation of Motion for the $\hat{\psi}$ Field
@@
$J_{1}$ $=v_{g}^{2}\Big{[}V+\Delta(1-x^{2})-\Delta^{2}x^{2}+2(1+\Delta)x(1-x^{2})\partial_{x}-(1-x^{2})^{2}\partial_{x}^{2}\Big{]},$
$J_{2}$ $=v_{g}\Big{[}iv_{g}+1-x^{2}-2\Delta x^{2}+2x(1-x^{2})\partial_{x}\Big{]}.$ (A9)
@@
$L_{1}^{a}$ $=v_{g}^{2}\Big{[}\tilde{V}-(1+\Delta)\Delta-2(1+\Delta)x\partial_{x}+(2-x^{2})\partial_{x}^{2}\Big{]},$ $L_{2}^{a}=v_{g}\Big{[}{-}1-2\Delta-2x\partial_{x}\Big{]},$
$J_{1}^{b}$ $=v_{g}^{2}\Big{[}V_{0}+xV_{1}+\Delta+2(1+\Delta)x\partial_{x}-\partial_{x}^{2}\Big{]},$ $J_{2}^{b}=v_{g}\Big{[}iv_{g}+1+2x\partial_{x}\Big{]},\hskip 8.0pt$ (A11)
@@
A.4 4. Determinant of the Matrix $L^{N}-\Omega\text{Id}$
@@
The QNM spectrum is given when this determinant vanishes, meaning we can scale $\Omega^{2}\text{Id}-i\Omega L_{2}^{N}-L_{1}^{N}$ without changing our result. First, we substitute for the discretized operators $L_{1}^{N}$ and $L_{2}^{N}$, making use of the asymptotic dispersion relation to rewrite $v_{g}$ and $v_{p}$ in terms of the real and imaginary parts of $\Omega$. Then, after an overall scaling of the matrix, we arrive at the condition $\det(M)=0$, with